# XAI-Based Detection of Adversarial Attacks on Deepfake Detectors

### v1

@@ 1 Introduction
Deepfake technology, which involves generating realistic media content, has advanced significantly in recent years, leading to the creation of increasingly sophisticated and convincing fake videos, images, and audio recordings (Westerlund, 2019; Khanjani et al., 2021; Guarnera et al., 2020b). In response, there has been a growing interest in developing deepfake detectors capable of identifying and flagging these manipulated media (Le & Woo, 2023; Hou et al., 2023; Rana et al., 2022; Lyu, 2020; Dolhansky et al., 2020; Guarnera et al., 2020a). However, as with any security system, these detectors are vulnerable to adversarial attacks, which aim to deceive or manipulate detector outputs by making subtle changes to the input, highlighting the need for robust detection methods.
@@ 2 Previous Work
Transitioning to adversarial attacks, the field of machine learning security confronts the persistent challenge posed by adversarial examples. These intentionally crafted inputs exploit vulnerabilities in deep neural network models across diverse domains, compelling erroneous predictions (Szegedy et al., 2014). Noteworthy are the gradient-based attacks, extensively explored in the literature, which exploit gradients to generate adversarial perturbations (Papernot et al., 2015; Lapid & Sipper, 2023a; Papernot et al., 2017; Li & Lyu, 2019; Moosavi-Dezfooli et al., 2017; Tamam et al., 2023; Carlini & Wagner, 2017a; Lapid & Sipper, 2023b; Eykholt et al., 2018; Shi et al., 2019; Lapid et al., 2022; Carlini & Wagner, 2018; Qin et al., 2019; Neekhara et al., 2019; Ebrahimi et al., 2018; Lapid et al., 2023; Belinkov & Bisk, 2018; Hagiwara et al., 2019).
@@ 3.1 Dataset Preparation
To evaluate the robustness of our methodology, we attack the FF+ dataset (see Section 4) using 4 different attacks: Projected Gradient Descent (PGD) (Madry et al., 2017), Auto Projected Gradient Descent (APGD) (Croce & Hein, 2020), Natural Evolution Strategies (NES) (Qiu et al., 2021) and Square Attack (Andriushchenko et al., 2020). The adversarial detector was trained using PGD only; the other attacks were used for testing only. For training our model the attack process iteratively applied PGD with a fixed maximum perturbation $\epsilon=16/255$ to generate manipulated instances, resulting in a dataset twice the original size. Subsequently, interpretability maps were generated across this augmented dataset using XAI techniques outlined in Section 3.4, facilitating the assessment of our approach’s sensitivity to manipulated features for discriminating between authentic and manipulated videos.
@@ 3.3 Adversarial Attacks
In this section we describe 4 attack methods that we use for creating the attacked-videos dataset. Two are white-box attacks: PGD (Madry et al., 2017) and Auto Projected Gradient Descent (APGD) (Croce & Hein, 2020). Two are black-box attacks: Natural Evolution Strategies (NES) (Qiu et al., 2021) and Square Attack (Andriushchenko et al., 2020).
@@
$\begin{split}\mathcal{L}&=\max(Z(x)_{\text{Real}}-Z(x)_{\text{Fake}},0).\end{split}$ (1)
Here, $Z(x)_{y}$ is the logit of label $y$, where $y\in\{\text{Real},\text{Fake}\}$. Minimizing the above loss function maximizes the score for our target label Real. The loss function adopted in this study, as advocated by Carlini & Wagner (2017b), was specifically chosen for its empirical effectiveness in generating adversarial samples with reduced distortion, and its demonstrated resilience to defensive distillation strategies. We used PGD to optimize the above objective while constraining the magnitude of the perturbation as follows:
@@ 3.5 Adversarial Attack Detection
For the creation of our datasets we employed both real and fake videos from the FF+ dataset, creating sets of real or attacked videos and XAI maps. We then formed pairs of real/attacked images and their corresponding XAI maps. Leveraging a pretrained backbone model, e.g., ResNet50 (He et al., 2016), we generated embeddings (feature vectors) for these images. To classify them into real or attacked categories, we applied two layers of linear transformation followed by an activation function. This composite architecture is herein referred to as the Detect-ResNet50 model.
@@
Figure 3: Frame analysis in suspected deepfake videos involves face extraction and classification using deepfake detectors. If classified as fake, the frame is labeled as fake. If classified as real, the face undergoes XAI map creation. The resulting XAI map and face are processed through a backbone model to generate embeddings, which are then input into the classification head to determine ‘real’ or ‘attacked’ status.
@@ 4 Experiments
For all attacks conducted we set a perturbation constraint of $\epsilon=16/255$. In the case of white-box attacks, the hyperparameters used were as follows: for the PGD attack—a maximum of $100$ iterations; for the APGD attack—a maximum of $100$ iterations and $5$ restarts. For black-box attacks the hyperparameters were: for the NES attack—a step size of $1/255$, a maximum of $100$ iterations, and $5$ restarts; for the Square attack—a maximum of $5000$ iterations, and the sampling-distribution parameter was set to $0.8$.
@@
Table 1 summarizes the pre-attack detection results, i.e., before performing adversarial attacks. Table 2 summarizes the post-attack detection results on deepfake images. The latter shows that attacked images cause the deepfake detector to lose its ability to ascertain real from not.
@@ 5 Results
Model Attack G-Backprop Inp $\times$ Grad Int-Grad Saliency
EfficientNetB4ST ($*{}^{*})$ (85.04%) (84.61%) (85.00%) (89.16%)
$*{}^{*}$ 99.60% 97.54% 94.48% 99.60%
@@
Total Average 85.95% 83.20% 84.07% 85.72%
Model Attack G-Backprop Inp $\times$ Grad Int-Grad Saliency
EfficientNetB4ST ($*{}^{*})$ (69.00%) (74.65%) (78.26%) (78.08%)
$*{}^{*}$ 93.46% 92.36% 91.58% 98.02%
@@
Total Average 83.45% 71.01% 75.40% 82.15%
For the finetuned ResNet50 model (Table 3), we observe that the average accuracy across all evaluated attacks ranges from 83.20% to 85.95%. Notably, the G-Backprop attack exhibits the highest success rate, achieving an average accuracy of 85.95%. This is closely followed by the Saliency attack, with an average accuracy of 85.72%. Conversely, the Inp $\times$ Grad and Int-Grad attacks demonstrate relatively lower average accuracies of 83.20% and 84.07%, respectively.
Performance slightly declined when confronted with the Square attack, while using EfficientNetB4ST, with accuracies ranging from 61.36% to 63.13%. This decrease in performance can be attributed to the model’s training using PGD only, indicating a potential limitation in generalization to other adversarial attack types. On the other hand, the performance of our approach against Square Attack on XceptionNet is better, reaching a best score of 82.20% with Saliency.
@@
Turning our attention to the second configuration of the ResNet50 model (Table 4), i.e., finetuning classification head only, we note similar trends in attack effectiveness. Across all evaluated attacks, the average accuracy ranges from 71.01% to 83.45%. Once again, the G-Backprop attack emerges as the most successful, achieving an average accuracy of 83.45%. This is followed by the Saliency attack, with an average accuracy of 82.15%. Conversely, the Inp $\times$ Grad and Int-Grad attacks exhibit lower average accuracies of 71.01% and 75.40%, respectively.
@@
Our findings underscore the balance between finetuning the entire model versus specific layers, adding another consideration. Finetuning the entire model allows for thorough adaptation to the target domain, enhancing task alignment and performance against adversarial attacks. Yet, it requires significant computational resources and may not be viable in resource-limited settings. On the other hand, finetuning only certain layers provides a more practical option, reducing computational burden while enabling some domain adaptation. However, it restricts model adaptation and may result in lower performance, especially against advanced adversarial attacks.
@@
Overall, we have assessed the performance of different XAI methods, with varying degrees of success, depending on the attack. Our findings emphasize the importance of developing robust defense strategies to safeguard against the proliferation of synthetic manipulations in digital media.
@@ 6 Discussion
Effectiveness of XAI in adversarial detection. The use of XAI techniques in the context of detecting adversarial attacks on deepfake detectors has demonstrated significant promise. The interpretability provided by XAI methods, such as feature attribution and saliency maps, enhances our understanding of model decision-making processes. In our experiments the integration of XAI facilitated the identification of subtle adversarial manipulations that might have otherwise gone unnoticed, as shown in the black XAI image experiments (Table 3 and Table 4). This highlights the importance of incorporating explainability mechanisms in deepfake detection systems to improve their robustness against adversarial attacks.
@@
Limitations. While our proposed adversarial detector for attacks on deepfake detectors demonstrates promising results, its efficacy may be limited by the diversity of adversarial attacks it has been trained on, and its generalization capability across different models and datasets. Moreover, the possibility of adversaries targeting our detector itself and the ethical implications of its deployment underscore the need for ongoing research to address these challenges and ensure its practical applicability and societal impact. However, it is reasonable to assume that the attacker lacks access to the adversarial detector.

### v2

@@ Abstract
22footnotetext: Code available at https://github.com/razla/XAI-Based-Detection-of-Adversarial-Attacks-on-Deepfake-Detectors33footnotetext: This research was supported by the Israeli Innovation Authority through the Trust.AI consortium
@@ 1 Introduction
Deepfake technology, which involves generating realistic media content, has advanced significantly in recent years, leading to the creation of increasingly sophisticated and convincing fake videos, images, and audio recordings (Westerlund, 2019; Khanjani et al., 2021; Guarnera et al., 2020b). In response, there has been a growing interest in developing deepfake detectors capable of identifying and flagging these manipulated media (Le & Woo, 2023; Hou et al., 2023; Lapid et al., 2024a; Rana et al., 2022; Lyu, 2020; Dolhansky et al., 2020; Guarnera et al., 2020a). However, as with any security system, these detectors are vulnerable to adversarial attacks, which aim to deceive or manipulate detector outputs by making subtle changes to the input, highlighting the need for robust detection methods.
@@ 2 Previous Work
Transitioning to adversarial attacks, the field of machine learning security confronts the persistent challenge posed by adversarial examples. These intentionally crafted inputs exploit vulnerabilities in deep neural network models across diverse domains, compelling erroneous predictions (Szegedy et al., 2014). Noteworthy are the gradient-based attacks, extensively explored in the literature, which exploit gradients to generate adversarial perturbations (Papernot et al., 2015; Lapid & Sipper, 2023a; Papernot et al., 2017; Li & Lyu, 2019; Moosavi-Dezfooli et al., 2017; Tamam et al., 2023; Carlini & Wagner, 2017a; Lapid & Sipper, 2023b; Eykholt et al., 2018; Shi et al., 2019; Lapid et al., 2022; Carlini & Wagner, 2018; Qin et al., 2019; Neekhara et al., 2019; Ebrahimi et al., 2018; Lapid et al., 2024b; Belinkov & Bisk, 2018; Hagiwara et al., 2019).
@@ 3.1 Dataset Preparation
To evaluate the robustness of our methodology, we attack the FF+ dataset (see Section 4) using 4 different attacks: Projected Gradient Descent (PGD) (Madry et al., 2017), Fast Gradient Sign Method (Goodfellow et al., 2014), Auto Projected Gradient Descent (APGD) (Croce & Hein, 2020), Natural Evolution Strategies (NES) (Qiu et al., 2021), and Square Attack (Andriushchenko et al., 2020). Herein we focus on the $\|\cdot\|_{\infty}$ norm constraint. The adversarial detector was trained using PGD only; the other attacks were used for testing only. For training our model the attack process iteratively applied PGD with a fixed maximum perturbation $\epsilon=16/255$ to generate manipulated instances, resulting in a dataset twice the original size. Subsequently, interpretability maps were generated across this augmented dataset using XAI techniques outlined in Section 3.4, facilitating the assessment of our approach’s sensitivity to manipulated features for discriminating between authentic and manipulated videos.
@@ 3.3 Adversarial Attacks
In this section we describe 5 attack methods that we use for creating the attacked-videos dataset. Three are white-box attacks: PGD (Madry et al., 2017), FGSM (Goodfellow et al., 2014), and Auto Projected Gradient Descent (APGD) (Croce & Hein, 2020). Two are black-box attacks: Natural Evolution Strategies (NES) (Qiu et al., 2021) and Square Attack (Andriushchenko et al., 2020).
@@
$\begin{split}\mathcal{L}&=\max(Z(x)_{\texttt{real}}-Z(x)_{\texttt{fake}},0).\end{split}$ (1)
Here, $Z(x)_{y}$ is the logit of label $y$, where $y\in\{\texttt{real},\texttt{fake}\}$. Minimizing the above loss function maximizes the score for our target label Real. The loss function adopted in this study, as advocated by Carlini & Wagner (2017b), was specifically chosen for its empirical effectiveness in generating adversarial samples with reduced distortion, and its demonstrated resilience to defensive distillation strategies. We used PGD to optimize the above objective while constraining the magnitude of the perturbation as follows:
@@
FGSM. The FGSM method, pioneered by Goodfellow et al. (2014), operates by perturbing input data based on the sign of the gradient of the loss function with respect to the input. This perturbation is scaled by a small constant, $\epsilon$, to ensure that changes are imperceptible but effective in fooling the deepfake detector. FGSM is simple and efficient, making it a popular choice for crafting adversarial examples. Despite its simplicity, FGSM can achieve significant perturbations, underscoring its effectiveness in bypassing deepfake detection systems.
@@ 3.5 Adversarial Attack Detection
For the creation of our datasets we employed both real and fake videos from the FF+ dataset, creating sets of real or attacked videos and XAI maps. We then formed pairs of unattacked/attacked images and their corresponding XAI maps. Leveraging a pretrained backbone model, e.g., ResNet50 (He et al., 2016), we generated embeddings (feature vectors) for these images. To classify them into unattacked or attacked categories, we applied two layers of linear transformation followed by an activation function. This composite architecture is herein referred to as the Detect-ResNet50 model.
@@
Figure 3: Frame analysis in suspected deepfake videos involves face extraction and classification using deepfake detectors. If classified as fake, the frame is labeled as fake. If classified as real, the face undergoes XAI map creation. The resulting XAI map and face are processed through a backbone model to generate embeddings, which are then input into the classification head to determine ‘unattacked’ or ‘attacked’ status.
Note that we now have two “axes” of interest, as it were: real vs. fake images, and attacked vs. unattacked images.
@@ 4 Experiments
Herein we focus on the $\|\cdot\|_{\infty}$ norm constraint. For all attacks conducted we use a perturbation constraint of $\epsilon=16/255$. In the case of white-box attacks, the hyperparameters used were as follows: for the PGD attack—a maximum of $100$ iterations; for FGSM—default configuration (1 iteration); for the APGD attack—a maximum of $100$ iterations and $5$ restarts. For black-box attacks the hyperparameters were: for the NES attack—a step size of $1/255$, a maximum of $100$ iterations, and $5$ restarts; for the Square attack—a maximum of $5000$ iterations, and the sampling-distribution parameter was set to $0.8$.
@@
Table 1 summarizes the pre-attack detection results, i.e., before performing adversarial attacks.
@@
Table 2 summarizes the post-attack detection results on deepfake images. The latter shows that attacked images cause the deepfake detector to lose its ability to ascertain real from not.
@@
In addition, our study incorporates adaptive attack strategies into the evaluation framework. Importantly, our proposed detection methodology assumes the adversary does not possesses simultaneous access both to deepfake detection and to adversarial detection mechanisms. We believe this assumption accords with real-life scenarios—such as those encountered in social media platforms, where an attacker’s actions are limited to uploading a video without knowledge of the underlying modules operating within the application. We advocate for the deployment of our approach in real-world settings, emphasizing its accessibility exclusively through API calls. Herein, we used a simple loss function:
$\mathcal{L}_{\text{standard-adaptive}}=\mathcal{L}_{\text{BCE}}(f(x+\delta),\texttt{real})+\mathcal{L}_{\text{BCE}}(g(x+\delta),\texttt{unattacked}),$ (3)
where BCE is the Binary Cross-Entropy Loss, $f$ is the deepfake detector, real is the ‘real’ class, $g$ is the adversarial detector, and ben is the ‘benign’ class. Thus the adversary’s goal is to minimize $\mathcal{L}_{\text{standard-adaptive}}$.
Moreover, we added another loss function, which tries to minimize the distance of the resultant XAI map:
$\mathcal{L}_{\text{XAI-adaptive}}=\mathcal{L}_{\text{BCE}}(f(x+\delta),\texttt{real})+\|(\text{XAI}(x+\delta;f),\text{XAI}(x;f))\|_{2},$ (4)
where XAI is a function that takes an input and a model, and outputs a XAI map. Again, the adversary’s goal is to minimize this loss function.
Finally, we incorporated generalization assessments into our study, wherein we evaluated the efficacy of an adversarial detection model initially trained on XceptionNet against the alternative architecture of EfficientNetB4ST, and, conversely, we examined the performance of the model trained on EfficientNetB4ST when confronted with XceptionNet.
4.1 Computational Overhead Experiments
To comprehensively assess the computational overhead introduced by integrating XAI techniques, we designed an experiment specifically focusing on measuring the additional computational costs. The experimental setup included a high-performance computing environment with an Intel Xeon Gold 6136 @ 3Ghz CPU, NVIDIA GRID P40-24Q, and 64GB of RAM. The software environment consisted of Windows 10, Python 3.9, Pytorch 2.0.1, and the Captum library for implementing XAI techniques (Kokhlikyan et al., 2019) . We measured the processing time and resource usage (CPU/GPU utilization and memory consumption) of our adversarial detection model, both with and without the integrated XAI methods.
The experiment was structured as follows:
  • Baseline Measurement: Run the adversarial-detection model without any XAI techniques, recording the processing time and memory consumption.
  • XAI Integration Measurement: Integrate the selected XAI techniques into the pipeline and repeat the measurements.
  • Comparison: Compare the baseline and XAI integration measurements to quantify the overhead.
@@ 5 Results
Deepfake Detector Attack G-Backprop Inp $\times$ Grad Int-Grad Saliency
EfficientNetB4ST (${}^{*})$ (85.04%) (84.61%) (85.00%) (89.16%)
PGD 99.60% 97.54% 94.48% 99.60%
FGSM 91.90% 89.93% 85.62% 93.11%
@@
Total Average 83.03% 82.92% 82.33% 87.46%
Deepfake Detector Attack G-Backprop Inp $\times$ Grad Int-Grad Saliency
EfficientNetB4ST (${}^{*})$ (69.00%) (74.65%) (78.26%) (78.08%)
PGD 93.46% 92.36% 91.58% 98.02%
FGSM 76.65% 82.26% 83.70% 88.02%
@@
Total Average 77.54% 75.96% 75.21% 81.85%
For the finetuned ResNet50 model (Table 3), we observe that the average accuracy across all evaluated attacks ranges from 75.76% to 90.30%. Notably, Saliency exhibits the highest success rate at detecting the attacks, achieving an average accuracy of 87.46%. Conversely, G-Backprop, Inp $\times$ Grad and Int-Grad demonstrate relatively lower average accuracies of 83.03%, 82.92% and 82.33%, respectively.
Performance slightly declined when confronted with the Square attack, while using EfficientNetB4ST, with accuracies ranging from 61.36% to 63.13% across all the XAI techniques. This decrease in performance can be attributed to the model’s training using PGD only, indicating a potential limitation in generalization to other adversarial attack types. On the other hand, the performance of our approach against Square Attack on XceptionNet is better, reaching a best score of 82.20% with Saliency.
@@
Turning our attention to the second configuration of the ResNet50 model (Table 4), i.e., finetuning the classification head only, we note similar trends in attack effectiveness. Across all evaluated attacks, the average accuracy ranges from 67.00% to 89.41%. Once again, Saliency emerges as the most successful, achieving an average accuracy of 81.85%. This is followed by G-Backprop, with an average accuracy of 77.54%. Conversely, Inp $\times$ Grad and Int-Grad exhibit lower average accuracies of 75.96% and 75.21%, respectively.
@@
Our findings underscore the balance between finetuning the entire model versus specific layers, adding another consideration. Finetuning the entire model allows for thorough adaptation to the target domain, enhancing task alignment and performance against adversarial attacks. Yet, it requires significant computational resources and may not be viable in resource-limited settings. On the other hand, finetuning only certain layers provides a more practical option, reducing computational burden while enabling some domain adaptation. However, it restricts model adaptation and may result in lower performance, especially against advanced adversarial attacks. On average, the detection rate on EfficientNetB4ST is better than the detection rate on XceptionNet.
Table 7 and Table 7 show results of the adaptive attacks, which are aware of both the deepfake detector and the adversarial detector, while using Equation 3 and Equation 4 respectively. We emphasize again that we do not think it is a realistic scenario and our detector was not trained against these kinds of attacks. In the first scenario, using Equation 3, the attacks successfully fooled both detectors, with EfficientNetB4ST showsing more resillience than XceptionNet. Using Equation 4, EfficientNetB4ST shows excellent performance, while XceptionNet still fails to resist these attacks.
Table 5: Comparison of Adaptive Attacks
Table 6: Adaptive attacks using Equation 3.
XceptionNet EfficientNetB4ST
Vid F2F Vid F2F
PGD 0.00% 0.00% 0.00% 0.16%
FGSM 0.00% 0.49% 80.00% 88.50%
Table 7: Adaptive attacks using Equation 4.
XceptionNet EfficientNetB4ST
Vid F2F Vid F2F
PGD 0.00% 0.24% 100.00% 98.60%
FGSM 0.00% 0.25% 100.00% 98.30%
In addition we conducted a generalization study where we evaluated the generalization of our approach: training on one model and evaluating on another. The results are delineated in Table 8, where we see generalization performance ranging from 66.00% to 82.97%. We note that on average training on XceptionNet’s attacked images and evaluating on EfficientNetB4ST yields the best results, ranging from 75.05%-82.97%.
Table 8: Transferability results for various attacks. A pretrained adversarial detector trained on a backbone deepfake detector is tested on attacks that were optimized on another backbone detector.
Source Model $\rightarrow$ Transferred Model Attack G-Backprop Inp $\times$ Grad Int-Grad Saliency
EfficientNetB4ST $\rightarrow$ XceptionNet PGD 67.33% 68.03% 80.50% 71.68%
FGSM 67.38% 68.02% 80.67% 71.67%
APGD 70.30% 63.37% 82.80% 85.30%
Square 47.10% 66.56% 82.93% 76.45%
NES 77.90% 68.08% 81.64% 99.60%
[…]
XceptionNet $\rightarrow$ EfficientNetB4ST PGD 95.90% 86.84% 83.21% 90.68%
FGSM 50.88% 73.67% 65.22% 85.73%
APGD 89.80% 54.83% 77.85% 66.10%
Square 52.41% 81.98% 80.73% 82.18%
NES 86.35% 87.25% 83.32% 90.16%
[…]
Total Average 70.54% 71.86% 79.88% 81.95%
@@
Overall, we have assessed the performance of different XAI methods, with varying degrees of success, depending on the attack. Training an adversarial detector on EfficientNetB4ST yields a more-resilient detector than XceptionNet, on average. Our findings emphasize the importance of developing robust defense strategies to safeguard against the proliferation of synthetic manipulations of digital media.
5.1 Computational Overhead Results
The results of our computational overhead experiments are presented in Table 9, which summarizes the processing-times metrics for both the baseline model and the pipeline with integrated XAI techniques.
Table 9: Computational overhead of our proposed pipeline.
Model Average Processing Time (ms)
Baseline G-Backprop Inp $\times$ Grad Int-Grad Saliency
XceptionNet 222.0 338.1 (+52%) 269.4 (+21%) 3640.3 (+1539%) 264.6 (+19%)
EfficientNet 247.1 361.5 (+46%) 335.0 (+36%) 8440.3 (+3316%) 358.1 (+45%)
These results indicate that while XAI techniques enhance explainability, they also introduce varying degrees of computational overhead. The Integrated Gradients method, in particular, incurs substantial overhead, making it less practical for real-time applications, compared to other XAI techniques. Conversely, methods like G-Backprop and Input $\times$ Gradient offer a more balanced trade-off between computational cost and explainability, exhibiting relatively lower processing times while still providing valuable insights into model decisions.
In addition to processing time, we also measured the memory overhead introduced by the XAI techniques. The baseline models have a memory usage of approximately 1.1GB. Integrating the XAI methods increases the memory usage by approximately 1GB, resulting in a total memory consumption of around 2.1GB. This additional memory overhead is primarily due to the storage requirements for the gradients and intermediate computations used by the XAI techniques.
The increased memory footprint is a crucial consideration for deploying these models in real-world settings, especially on resource-constrained devices. While the overhead is significant, it remains manageable within the context of modern computing environments equipped with sufficient memory resources. This analysis underscores the importance of balancing the need for explainability with the available computational resources, particularly in applications where memory constraints are a critical factor.
@@ 6 Discussion
Effectiveness of XAI in adversarial detection. The use of XAI techniques in the context of detecting adversarial attacks on deepfake detectors has demonstrated significant promise. The interpretability provided by XAI methods, such as feature attribution and saliency maps, enhances our understanding of model decision-making processes. In our experiments the integration of XAI facilitated the identification of subtle adversarial manipulations that might have otherwise gone unnoticed, as shown in the black XAI image experiments (Table 3 and Table 4). This highlights the importance of incorporating explainability mechanisms in deepfake detection systems to improve their robustness against adversarial attacks. Additionally, our findings reveal nuanced insights into the effectiveness of XAI techniques in mitigating adversarial attacks. Specifically, we observed that Saliency and G-Backprop outperform other methods in scenarios where the entire network is fine-tuned, while Saliency demonstrates superior performance when only the classification head is fine-tuned. Furthermore, on average, using EfficientNetB4ST’s XAIs yield better results than XceptionNet. This distinction underscores the importance of considering the architectural and training constraints when integrating XAI methods for robust deepfake detection.
Ethical Use of XAI. While this paper emphasizes the benefits of using XAI to enhance deepfake detectors, it is crucial to consider potential ethical concerns. XAI could potentially be misused by adversaries to reverse-engineer AI systems and exploit vulnerabilities, thereby compromising the very robustness these techniques aim to enhance. To mitigate such risks, it is essential to implement strict security protocols governing the dissemination and application of XAI insights. This includes controlling access to sensitive interpretability data, collaborating with cybersecurity experts, and developing XAI methods that balance transparency with security. Continuous ethical deliberation within the research community is also necessary to ensure responsible AI usage.
Societal Implications. Enhanced deepfake detectors have significant societal benefits, such as preventing misinformation and protecting digital content integrity. However, their dual-use nature raises concerns about privacy and surveillance. These technologies could be misused to monitor individuals’ digital expressions, infringing on privacy rights. To address these concerns, comprehensive guidelines for the responsible deployment of deepfake detectors must be proposed and adhered to. These guidelines should ensure transparency, legal and ethical consistency, and protection of privacy rights. Multi-stakeholder dialogue involving policymakers, technologists, civil society, and the public is essential to define and enforce ethical boundaries for deploying these technologies.
@@
Limitations. While our proposed adversarial detector for attacks on deepfake detectors demonstrates promising results, its efficacy may be limited by the diversity of adversarial attacks it has been trained on, and its generalization capability across different models and datasets. Additionally, exploring a wider range of $\epsilon$ values in crafting adversarial examples could provide deeper insights into the robustness of the detector. Moreover, the possibility of adversaries targeting our detector itself and the ethical implications of its deployment underscore the need for ongoing research to address these challenges and ensure its practical applicability and societal impact. However, it is reasonable to assume that the attacker lacks access to the adversarial detector.